Solve the differential equation given by

(3x + 2y + y^2)dx + (x+4xy+5y^2)dy=0



1
Expert's answer
2021-08-25T13:53:24-0400
"\dfrac{\partial Q}{\partial x}=1+4y, \dfrac{\partial P}{\partial y}=2+2y"

"\dfrac{\partial Q}{\partial x}\not=\dfrac{\partial P}{\partial y}"

"\mu=x+y^2"

"(x+y^2)(3x + 2y + y^2)dx"

"+ (x+y^2)(x+4xy+5y^2)dy=0"

"P=3x^2+2xy+xy^2+3xy^2+2y^3+y^4"

"\dfrac{\partial P}{\partial y}=2x+8xy+6y^2+4y^3"

"Q=x^2+4x^2y+5xy^2+xy^2+4xy^3+5y^4"

"\dfrac{\partial Q}{\partial x}=2x+8xy+6y^2+4y^3"

"\dfrac{\partial P}{\partial y}=\dfrac{\partial Q}{\partial x}"

"\dfrac{\partial u}{\partial x}=3x^2+2xy+4xy^2+2y^3+y^4"

Integrate


"u=x^3+x^2y+2x^2y^2+2xy^3+xy^4+\varphi(y)"

Then

"\dfrac{\partial u}{\partial y}=x^2+4x^2y+6xy^2+4xy^3+\varphi'(y)"

"=Q=x^2+4x^2y+6xy^2+4xy^3+5y^4"

"\varphi'(y)=5y^4"

"\varphi(y)=y^5+C"

The original differential equation has the following solution


"x^3+x^2y+2x^2y^2+2xy^3+xy^4-y^5=C"